# Impact of flavour coupling on $SO(10)$-inspired leptogenesis

### v1

@@ 2.3 $SO(10)$-inspired leptogenesis
  • We have then the three neutrino unknowns $\theta_{23}$, $\delta$ and $m_{1}$. Here we have been very conservative not to contaminate the predictions from SO10INLEP with the partial experimental information we have. We have uniformly randomly generated the atmospheric neutrino mixing angle in the range $\theta_{23}=[38^{\circ},52^{\circ}]$ and the neutrino oscillation $C\!P$ violating phase in the full range $\delta=[-\pi,\pi]$. Finally we have uniformly logarithmically generated the lightest neutrinos mass in the range $\log(m_{1}/{\rm eV})=[-4,0]$.
  • The Majorana phases have been simply randomly uniformly generated in the full range $[0,2\pi]$.
  • The 6 parameters in $V_{L}$ have been uniformly randomly generated. The three mixing angles have been capped to the CKM matrix values, following $SO(10)$-inspired conditions, explicitly we adopted: $\theta_{12}^{L}\leq 13^{\circ}\simeq\theta_{12}^{CKM}\equiv\theta_{c}$, $\theta_{23}^{L}\leq 2.4^{\circ}\simeq\theta_{23}^{CKM}$, $\theta_{13}^{L}\leq 0.2^{\circ}\simeq\theta_{13}^{CKM}$. In any case the results do not depend on a precise choices of these upper bounds [30]. Finally, for the parameter $\alpha_{2}$, we have used throughout the paper a kind of maximum standard value, $\alpha_{2}\leq 5$.
@@ 2.4 Successful leptogenesis condition and solutions
identifies an allowed region in the space of the nine low energy neutrino parameters. In the approximation $V_{L}=I$ this would correspond to an eight-dimensional hypersurface. Turning on the three angles and three phases in $V_{L}$, this hypersurface turns into a layer with some thickness. It is important that it does not fill the whole parameter space and this yields constraints and predictions. In order to visualise this region, we show in Fig. 1 the three-dimensional projection,
@@
Figure 1: Scatter plot of the solutions obtained imposing successful SO10INLEP neglecting flavour coupling (left panel) and accounting for flavour coupling (right panel). The three grey areas correspond to the excluded regions by the three upper bounds on the absolute neutrino mass scale: Eq. (2) from cosmological observations, Eq. (10) from $0\nu\beta\beta$ and Eq. (11) from tritium beta decay. The three planes in light blue simply help understanding the 3-dim shape. Colour code: tauonic, muonic and strong thermal solutions are, respectively, yellow, green and blue points.
on the 3-dim space $(\delta,\theta_{23},m_{1})$ of the three unknown neutrino parameters [29], of a scatter plot obtained for $\alpha_{2}=5$ in the 15-dim space of the parameters in $U$ and $V_{L}$ by generating about $2\times 10^{6}$ solutions satisfying successful $SO(10)$-inspired leptogenesis and all low energy neutrino experimental constraints in Eq. (6). [^10] The success rate is approximately $0.01\%$ for the parameter ranges we have adopted in the scan. One can immediately notice how solutions are found only for $m_{1}\gtrsim 1\,{\rm meV}$, so that there is a clear lower bound on the absolute neutrino mass scale. The three dark grey regions indicate the upper bounds on neutrino masses Eq. (2) from cosmological observations, Eq. (10) from $0\nu\beta\beta$ and Eq. (11) from tritium beta decay.
@@ 2.4.1 Three types of solutions
There are three types of solutions in SO10INLEP [20]. These correspond to three different regions in the parameter space that can be partly distinguished in the three dimensional projection scatter plot in Fig. 1. The different colours indicate the flavour giving the main contribution to the final asymmetry, specifically:
  • The yellow points indicate tauon-flavour dominated solutions characterised by $K_{1\tau}\lesssim 1$. One can distinguish two distinct types of tauonic solutions: $\tau_{A}$ and $\tau_{B}$ solutions:
      The $\tau_{A}$ solutions are characterised by strong washout at the production ($K_{2\tau}\gtrsim 10$). Moreover, one has $K_{1\mu}\gg 1$, so that the muon contribution is many orders of magnitude below the observed value. For current allowed experimental values of $\theta_{23}$ as in Eq. (6), they can only be realised for $0.9\,{\rm meV}\lesssim m_{1}\lesssim 30\,{\rm meV}$. The blue points indicate the subset of tauon flavour dominated solutions that realise ST-SOINLEP and that we will discuss in Section 5. The $\tau_{A}$ solutions can realise ST-SOINLEP while the $\tau_{B}$ cannot.
[…]
@@
It is also interesting to notice that for all $SO(10)$-inspired solutions one has $K_{2e}\ll K_{2\mu}\simeq K_{2\tau_{2}^{\bot}}$. This means that the electronic component of the leptons produced from $N_{2}$-decays is negligible. Essentially, all leptons produced from $N_{2}$ decays have a flavour composition that lies quite precisely on the muon-tauon plane. The reason can be easily understood analytically. Indeed from Eq. (31) and Eq, (17) one easily derives $K_{2e}/K_{2\mu}\lesssim|V_{L21}|^{2}\sim\theta^{2}_{\rm c}\sim 0.05$. One can see from Eq. (27) that this result, combined with the fact that also $\varepsilon_{2e}\ll\varepsilon_{2\mu}$, implies that phantom terms are always negligible.
In Fig. 2 we plot RH neutrino masses, decay parameters and asymmetries versus $m_{1}$ for three benchmark points in the parameter space and for $(\alpha_{1},\alpha_{2},\alpha_{3})=(1,5,1)$ (the values of $\alpha_{1}$ and $\alpha_{3}$ only affect $M_{1}$ and $M_{3}$). For all of them one has $\eta_{B}^{\rm lep}\geq\eta_{B}^{\rm exp}$ within some range of $m_{1}$ values. Therefore, successful SO10INLEP is realised for some special values of $m_{1}$ where $\eta_{B}^{\rm lep}=\eta_{B}^{\rm exp}$. Each of these three benchmark points provides an example of solution belonging to one of the three types: $\tau_{A}$ (left panels), $\tau_{B}$ (central panels) and $\mu$ (right panels). At this stage of the discussion, one should focus only on the the dashed lines, corresponding to neglect flavour coupling.
@@
The values of the different parameters for each solution are given in Table 1. The muonic solution has been selected among those with a sizeable sub-dominant tauonic contribution to the asymmetry. This can be clearly seen in the bottom right panel (yellow dashed-line), the tauonic contribution to the total asymmetry is $\sim 6\%$. However, it requires a value $m_{1}=45\,{\rm meV}$ in tension with the current cosmological upper bound.
@@
$\tau_{A}$ $33.44^{\circ}$ $8.61^{\circ}$ $43.12^{\circ}$ $-62.4^{\circ}$ 0.31 0.92 $9.46^{\circ}$ $0.16^{\circ}$ $0.0445^{\circ}$ 0.35 0.49 1.59
$\tau_{B}$ $30.9^{\circ}$ $8.2^{\circ}$ $41.5^{\circ}$ $-34.59^{\circ}$ 0.952 0.920 $9.90^{\circ}$ $0.035^{\circ}$ $0.86^{\circ}$ 0.66 0.99 1.99
@@ 2.4.2 Lower bound on $m_{1}$
The major feature that can be noticed in the three-dimensional scatter plot in Fig. 1 is the existence of a lower bound on $m_{1}$. It is also clear that this lower bound depends both on $\theta_{23}$ and even more strongly on $\delta$. In the light of an accurate measurement of these two neutrino mixing unknowns at DUNE+T2HK, it is then important to derive the value of the lower bound on $m_{1}$ for each point in the plane $\theta_{23}$-$\delta$ In Fig. 3 we show iso-contour lines of the lower bound on $m_{1}$ in the plane $\delta$ versus $\theta_{23}$.
Figure 3: Values of the lower bound on $m_{1}$ in the $\theta_{23}-\delta$ plane (isocontour lines) without flavour coupling (left panel) and with flavour coupling. The white dashed lines are the $1\sigma$, $2\sigma$and $3\sigma$ experimental constraints from [23]. The orange area is the area excluded by the cosmological upper bound (2). The green area is the portion of the plane that opens up thanks to the muonic solutions from flavour coupling effects.
We also superimpose the two-dimensional experimental constraints from latest global analyses [23]. It can be seen how at $3\sigma$ the global lower bound is $m_{1}\gtrsim 0.9\,{\rm meV}$. An interesting feature is that there is a large region of the plane $\theta_{23}$-$\delta$, in red colour, that is incompatible with the cosmological upper bound Eq. (1) and that the $1\sigma$ experimental regions fall almost entirely just within this region. This clearly shows how SO10INLEP is being strongly tested by low energy neutrino experimental results. Therefore, it is very important to have a clear understanding of theoretical uncertainties to draw firm conclusions from a comparison between theoretical calculations and experimental constraints. As we will see in the next section, flavour coupling plays an important role in this respect.
@@ 2.4.3 Upper bound on $\theta_{23}$
Figure 4: Two-dimensional projection on the plane $m_{1}$-$\theta_{23}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel).
[…]
@@ 2.4.4 Excluded regions for $\delta$
Figure 5: Two-dimensional projection on the plane $m_{1}$-$\theta_{23}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel). As in Fig. 1, the vertical grey areas denote the excluded regions by the three upper bounds on $m_{1}$: Eq. (2) from cosmological observations, Eq. (10) from $0\nu\beta\beta$ and Eq. (11) from tritium beta decay. The horizontal grey area is the $3\sigma$ excluded $\theta_{2}3$ range of values from Eq. (6).
[…]
@@ 2.4.5 Majorana phases
In the left panel of Fig 6 we show the projection of the scatter plot solutions on the $\sigma-\rho$ plane. One can notice how large part of the plane is excluded. This is because, as well understood analytically [12, 33], the Majorana phases need to be within certain ranges in order for either $K_{1\tau}\lesssim 1$ or $K_{1\mu}\lesssim 1$ (or both).
@@
Figure 6: Two-dimensional projection on the plane $\sigma$-$\rho$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel).
Notice how in this plane the muon solution allowed regions (green points) are well disconnected by the tauon solution allowed region (yellow points) and lay around the points $(\sigma,\rho)=((2n+1)\pi/2,m\pi)$, with $m,n\in{\mathbb{Z}}$. These are indeed the values that make possible to have $K_{1\mu}\lesssim 1$ [33].
In Fig. 7 we also shown the projection on the plane $m_{1}-\sigma$. It can be clearly seen how the values of $\sigma$ for the muon solutions are different from those for the $\tau$ solutions.
@@
Figure 7: Two-dimensional projection on the plane $m_{1}$-$\theta_{23}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel).
@@ 2.4.6 $0\nu\beta\beta$ effective neutrino mass
The observation of a $0\nu\beta\beta$ signal would be a crucial discovery supporting, in general, leptogenesis since it would be a clear sign that lepton number is violated at tree level. Specifically, for SO10INLEP, it would provide the measurement of an additional low energy neutrino observable depending on all nine low energy neutrino parameters. This would represent another important experimental testing experimental constraint. In Fig. 8 we show the allowed region in the traditional plane $m_{1}$-$m_{ee}$.
@@
Figure 8: Projection in the plane $m_{1}-m_{ee}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel).
The vertical grey area is the conservative excluded region from KamLAND-Zen upper bound given in Eq. (10), while the dashed line indicates the most stringent upper bound also given in Eq. (10). The thin black lines indicate the general allowed region compatible with the experimental constraints (6) (i.e., no successful SO10INLEP condition is imposed). One can see how SO10INLEP strongly restricts this region. Therefore, the simultaneous experimental determination of $m_{1}$ and $m_{ee}$ would provide a very strong test of SO10INLEP.

### v2

@@ 2.3 $SO(10)$-inspired leptogenesis
  • We have then the three neutrino unknowns $\theta_{23}$, $\delta$ and $m_{1}$. Here we have been very conservative not to contaminate the predictions from SO10INLEP with the partial experimental information we have. We have uniformly randomly generated the atmospheric neutrino mixing angle in the range $\theta_{23}=[38^{\circ},52^{\circ}]$ and the neutrino oscillation CPC\!P violating phase in the full range $\delta=[-\pi,\pi]$. Finally we have uniformly logarithmically generated the lightest neutrinos mass in the range $\log(m_{1}/{\rm eV})=[-4,0]$.
  • The Majorana phases have been simply randomly uniformly generated in the full range $0,2\pi]$.
  • The 6 parameters in $V_{L}$ have been uniformly randomly generated. The three mixing angles have been capped to the CKM matrix values, following $SO(10)$-inspired conditions, explicitly we adopted: $\theta_{12}^{L}\leq 13^{\circ}\simeq\theta_{12}^{CKM}\equiv\theta_{c}$, $\theta_{23}^{L}\leq 2.4^{\circ}\simeq\theta_{23}^{CKM}$, $\theta_{13}^{L}\leq 0.2^{\circ}\simeq\theta_{13}^{CKM}$. In any case the results do not depend on a precise choices of these upper bounds ST . Finally, for the parameter $\alpha_{2}$, we have used throughout the paper a kind of maximum standard value, $\alpha_{2}\leq 5$.
@@ 2.4 Successful leptogenesis condition and solutions
identifies an allowed region in the space of the nine low energy neutrino parameters. In the approximation $V_{L}=I$ this would correspond to an eight-dimensional hypersurface. Turning on the three angles and three phases in $V_{L}$, this hypersurface turns into a hyperlayer with some thickness. It is important that it does not fill the whole parameter space and this yields constraints and predictions. In order to visualise this region, we show in Fig. 1 the three-dimensional projection,
@@
Figure 1: Scatter plot of the solutions obtained imposing successful SO10INLEP neglecting flavour coupling (left panel) and accounting for flavour coupling (right panel). The three grey areas correspond to the excluded regions by the three upper bounds on the absolute neutrino mass scale: Eq. (2) from cosmological observations, Eq. (10) from $0\nu\beta\beta$ and Eq. (11) from tritium beta decay. The three planes in light blue simply help understanding the 3-dim shape. Colour code: tauonic, muonic and strong thermal solutions are denoted by yellow, green and blue points, respectively.
on the 3-dim space $(\delta,\theta_{23},m_{1})$ of the three unknown neutrino parameters DiBari:2020plh , of a scatter plot obtained for $\alpha_{2}=5$ in the 15-dim space of the parameters in $U$ and $V_{L}$ by generating about $2\times 10^{6}$ solutions satisfying successful $SO(10)$-inspired leptogenesis and all low energy neutrino experimental constraints in Eq. (6). [^11] The success rate is approximately $0.01\%$ for the parameter ranges we have adopted in the scan. One can immediately notice how solutions are found only for $m_{1}\gtrsim 1\,{\rm meV}$, so that there is a clear lower bound on the absolute neutrino mass scale. The three dark grey regions indicate the upper bounds on neutrino masses in Eq. (2) from cosmological observations, Eq. (10) from $0\nu\beta\beta$ and Eq. (11) from tritium beta decay.
@@ 2.4.1 Three types of solutions
There are three types of solutions in SO10INLEP riotto2 . These correspond to three different regions in the parameter space that can be partly distinguished in the three dimensional projection scatter plot in Fig. 1. The different colours indicate the flavour giving the main contribution to the final asymmetry, specifically:
  • The yellow points indicate tauon-flavour dominated solutions characterised by $K_{1\tau}\lesssim 1$. One can distinguish two distinct types of tauonic solutions, $\tau_{A}$ and $\tau_{B}$ solutions: [^12] The two corresponding regions in the multi-parameter space are not disjoint but the border line, hybrid, solutions are quite marginal. This means that the border between the two regions is quite thin, even though the two-dimensional projections we show might be misleading in that respect.
      The $\tau_{A}$ solutions are characterised by strong washout at the production ($K_{2\tau}\gtrsim 10$). Moreover, one has $K_{1\mu}\gg 1$, so that the muon contribution is many orders of magnitude below the observed value. For current allowed experimental values of $\theta_{23}$ as in Eq. (6), they can only be realised for $0.9\,{\rm meV}\lesssim m_{1}\lesssim 30\,{\rm meV}$. The blue points indicate the subset of tauon flavour dominated solutions that realise ST-SO10INLEP and that we will discuss in Section 5. The $\tau_{A}$ solutions can realise ST-SO10INLEP while the $\tau_{B}$ cannot in general [^13] Only a few special ones with large $K_{1\mu}\sim 10$ can be found..
      The $\tau_{B}$ solutions are characterised by a mild washout at the production with $2\lesssim K_{2\tau}\lesssim 10$. In this case one can also have simultaneously $K_{1\mu}\lesssim 1$ and in any case $K_{1\mu}\lesssim 10$. However, one has $K_{2\tau_{2}^{\bot}}\gg 10$. The muonic contribution is now only 2-3 orders of magnitudes smaller than the observed value. They can only be realised for $m_{1}\simeq(20$$60)\,{\rm meV}$.
@@
It is also interesting to notice that for all $SO(10)$-inspired solutions one has $K_{2e}\ll K_{2\mu}\simeq K_{2\tau_{2}^{\bot}}$. This means that the electronic component of the leptons produced from $N_{2}$-decays is negligible. Essentially, all leptons produced from $N_{2}$ decays have a flavour composition that lies quite precisely on the muon-tauon plane. The reason can be easily understood analytically: from Eq. (31) and Eq. (17) one easily derives $K_{2e}/K_{2\mu}\lesssim|V_{L21}|^{2}\sim\theta^{2}_{\rm c}\sim 0.05$. One can see from Eq. (27) that this result, combined with the fact that also $\varepsilon_{2e}\ll\varepsilon_{2\mu}$, implies that phantom terms are always negligible.
In Fig. 2 we plot RH neutrino masses, decay parameters and asymmetries versus $m_{1}$ for three benchmark points in the parameter space and for $(\alpha_{1},\alpha_{2},\alpha_{3})=(1,5,1)$ (the values of $\alpha_{1}$ and $\alpha_{3}$ only affect $M_{1}$ and $M_{3}$). For all of them one has $\eta_{B}^{\rm lep}\geq\eta_{B}^{\rm exp}$ within some range of $m_{1}$ values. Therefore, successful SO10INLEP is realised for some special values of $m_{1}$ where $\eta_{B}^{\rm lep}=\eta_{B}^{\rm exp}$. Each of these three benchmark points provides an example of solution belonging to one of the three types: $\tau_{A}$ (left panels), $\tau_{B}$ (central panels) and $\mu$ (right panels). At this stage of the discussion, one should focus only on the the dashed lines, obtained neglecting flavour coupling.
@@
The values of the different parameters for each solution are given in Table 1. The benchmark muonic solution has been selected among those with a sizeable sub-dominant tauonic contribution to the asymmetry. This can be clearly seen in the bottom right panel (yellow dashed-line), the tauonic contribution to the total asymmetry is $\sim 6\%$. However, it requires a value $m_{1}=45\,{\rm meV}$ in tension with the current cosmological upper bound.
@@
$\tau_{A}$ $33.44^{\circ}$ $8.61^{\circ}$ $43.12^{\circ}$ $62.4^{\circ}$ 0.31 0.92 $9.46^{\circ}$ $16^{\circ} 0.0$$0445^{\circ}$ 0.35 0.49 1.59
$\tau_{B}$ $30.9^{\circ}$ $8.2^{\circ}$ $41.5^{\circ}$ $34.59^{\circ}$ 0.952 0.920 $9.90^{\circ}$ $0.035^{\circ}$ $0.86^{\circ}$ 0.66 0.99 1.99
@@ 2.4.2 Lower bound on $m_{1}$
The major feature that can be noticed in the three-dimensional scatter plot in Fig. 1 is the existence of a lower bound on $m_{1}$. It is also clear that this lower bound depends both on $\theta_{23}$ and even more strongly on $\delta$. In the light of an accurate measurement of these two neutrino mixing unknowns at DUNE+T2HK, it is then important to derive the value of the lower bound on $m_{1}$ for each point in the plane $\theta_{23}$-$\delta$. In Fig. 3 we show iso-contour lines of the lower bound on $m_{1}$ in the plane $\delta$ versus $\theta_{23}$.
[…]
We also superimpose the two-dimensional experimental constraints from latest global analyses nufit24 . It can be seen how at $3\sigma$ the global lower bound is $m_{1}\gtrsim 0.9\,{\rm meV}$. An interesting feature is that there is a large region of the plane $\theta_{23}$-$\delta$, in red colour, that is incompatible with the cosmological upper bound Eq. (1) and that the $1\sigma$ experimental regions fall almost entirely just within this region. This clearly shows how SO10INLEP is being strongly tested by low energy neutrino experimental results. Therefore, it is very important to have a clear understanding of theoretical uncertainties to draw firm conclusions from a comparison between theoretical calculations and experimental constraints. As we will see in the next section, flavour coupling plays an important role in this respect.
@@ 2.4.3 Upper bound on $\theta_{23}$
Figure 4: Two-dimensional projection on the plane $m_{1}$-$\theta_{23}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel). Colour code as in Fig. 1. In the left (right) panel the 3 (2) stars (triangles) denote the 3 (2) benchmark solutions in Table 1 (2).
The most interesting feature is that, in the range $m_{1}\sim(10$$30)\,{\rm meV}$, the second octant is not allowed. This is another important clear test for SO10INLEP and one should notice that it is independent of the value of $\delta$.
@@ 2.4.4 Excluded regions for $\delta$
Figure 5: Two-dimensional projection on the plane $m_{1}$-$\theta_{23}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling (left) and accounting for flavour coupling (right). As in Fig. 1, the vertical grey areas denote the excluded regions by the three upper bounds on $m_{1}$: Eq. (2) from cosmological observations, Eq. (10) from $0\nu\beta\beta$ and Eq. (11) from tritium beta decay. The horizontal grey area is the $3\sigma$ excluded $\theta_{23}$ range of values from Eq. (6). Same conventions as in Fig. 4.
One can notice how the lower bound on $m_{1}$ is much more stringent in the half $\delta=180^{\circ}\pm 90^{\circ}$ compared to the other half range $\delta\sim 0\pm 90^{\circ}$. In particular, the current $1\sigma$ experimental range of values for $\delta$ falls in the first half that is allowed only for $m_{1}\gtrsim 10\,{\rm meV}$. It is also curious how there is a kind of hole for $\delta\sim 30^{\circ}-120^{\circ}$ and $m_{1}\sim(5$$15)\,{\rm meV}$. This shows again how a future precise determination of $\delta$ will represent a crucial test for SO10INLEP.
@@ 2.4.5 Majorana phases
In the left panel of Fig 6 we show the projection of the scatter plot solutions on the $\sigma-\rho$ plane. One can notice how large part of the plane is excluded. This is because, as well understood analytically decrypting , full , the Majorana phases need to be within certain ranges in order for either $K_{1\tau}\lesssim 1$ or $K_{1\mu}\lesssim 1$ (or both).
@@
Figure 6: Two-dimensional projection on the plane $\sigma$-$\rho$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel). Same conventions as in Fig. 4.
Notice how in this plane the muon solution allowed regions (green points) are well disconnected by the tauon solution allowed region (yellow points) and lay around the points $(\sigma,\rho)=((2n+1)\pi/2,m\pi)$, with $m,n\in{\mathbb{Z}}$. These are indeed the values that make possible to have $K_{1\mu}\lesssim 1$ full .
In Fig. 7 we also show the projection on the plane $m_{1}-\sigma$. It can be clearly seen how the values of $\sigma$ for the muon solutions are different from those for the $\tau$ solutions.
@@
Figure 7: Two-dimensional projection on the plane $m_{1}$-$\theta_{23}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling effects (right panel). Same conventions as in Fig. 4.
@@ 2.4.6 $0\nu\beta\beta$ effective neutrino mass
The observation of a $0\nu\beta\beta$ signal would be a crucial discovery supporting, in general, leptogenesis since it would be a clear sign that lepton number is violated at tree level. Specifically, for SO10INLEP, it would provide the measurement of an additional low energy neutrino observable depending on all nine low energy neutrino parameters. This would provide another important experimental constraint to be satisfied. In Fig. 8 we show the allowed region in the usual plane $m_{ee}$ versus $m_{1}$.
@@
Figure 8: Projection in the plane $m_{1}-m_{ee}$ of the scatter plot of the solutions obtained imposing successful SO10INLEP, neglecting flavour coupling effects (left panel) and accounting for flavour coupling (right panel). Same conventions as in Fig. 4. The grey lines delimit the general allowed region obtained from the experimental ranges in Eq. (6).
The vertical grey area is the conservative excluded region from the KamLAND-Zen upper bound in Eq. (10), while the dashed line indicates the most stringent upper bound also given in Eq. (10). The grey lines indicate the general allowed region compatible with the experimental constraints (6) (i.e., no successful SO10INLEP condition is imposed). One can see how SO10INLEP strongly restricts this region and, importantly, now there is a lower bound on $m_{ee}$ also in the range $2\,{\rm meV}\lesssim m_{1}\lesssim 7\,{\rm meV}$, even though we are considering NO. Therefore, the simultaneous experimental determination of $m_{1}$ and $m_{ee}$, realistically for both quantities $\gtrsim 1\,{\rm meV}$, would provide a very strong test in favour of SO10INLEP. This is true especially if $m_{1}$ is found in that particular range, also considering that the bulk of solutions, of $\tau_{A}$-type, do fall in that range. [^14] This argument contains some subjectivity depending whether one considers $m_{1}$ linearly or logarithmically: in the latter case, without SO10INLEP, one would expect not to measure $m_{ee}$ for $2\,{\rm meV}\lesssim m_{1}\lesssim 7\,{\rm meV}$. This expectation neglects that arbitrarily small values of $m_{ee}$ are obtained for fine-tuned phase cancellations. This also contributes to the level of subjectiveness. In any case, it is true, that with SO10INLEP such fine-tuned phase cancellations are not possible, since Majorana phases, as we have seen, are constrained within regions allowing $K_{1\tau}\lesssim 1$ and this justifies our statement that a simultaneous determination of $m_{1}$ and $m_{ee}$ would be a very strong test in favour of SO10INLEP.